# Unsupervised Meta-Learning via In-Context Learning

### v1

@@ Abstract
Unsupervised meta-learning aims to learn feature representations from unsupervised datasets that can transfer to downstream tasks with limited labeled data. In this paper, we propose a novel approach to unsupervised meta-learning that leverages the generalization abilities of in-context learning observed in transformer architectures. Our method reframes meta-learning as a sequence modeling problem, enabling the transformer encoder to learn task context from support images and utilize it to predict query images. At the core of our approach lies the creation of diverse tasks generated using a combination of data augmentations and a mixing strategy that challenges the model during training while fostering generalization to unseen tasks at test time. Experimental results on benchmark datasets, including miniImageNet, CIFAR-fs, CUB, and Aircraft, showcase the superiority of our approach over existing unsupervised meta-learning baselines, establishing it as the new state-of-the-art in the field. Remarkably, our method achieves competitive results with supervised and self-supervised approaches, underscoring the efficacy of the model in leveraging generalization over memorization.
@@ 1 Introduction
Various approaches have been proposed to address the UML problem [18, 19, 20, 23, 24, 25]. Despite its potential, UML faces several challenges. Existing UML methods often rely on simple data augmentations to construct the training tasks, while following the standard meta-learning task sampling pipeline for the evaluation phase. This results in a significant difference between training and testing tasks, constraining the model’s ability to generalize to unseen tasks during inference and requiring additional fine-tuning on the test domain. Furthermore, existing UML approaches typically assume that the training and test datasets belong to the same domain. In this work, we loosen this assumption resulting in a more challenging setting that necessitates better model generalization compared to usual meta-learning applications. We refer to this as the cross-domain scenario.
@@
In this paper, we propose a novel approach to UML that addresses these challenges by leveraging in-context learning within a transformer architecture [11, 31]. In-context learning [11] allows the model to use the context provided by a sequence of input-output pairs to make predictions about new inputs. Inspired by recent advancements in large language models (LLMs) [52, 4, 29], we formulate meta-learning as a sequence modeling problem, where a task is seen as a sequence of support images and an unknown query image. The support set is treated as the context utilized by the model to predict the class of the query image. We call our approach CAMeLU, which stands for Context-Aware Meta-Learning in Unsupervised scenarios. Central to our approach is a novel task creation mechanism that enables the generation of a large number of different tasks from an unlabeled dataset. Drawing inspiration from the natural decision process of learning by analogy [53], we construct tasks that closely resemble the structure of those encountered during inference. Specifically, we use a combination of different data augmentation techniques based on basic image manipulations [39] for generating the samples in the support set. Conversely, a strategy similar to mixup [57] is employed to generate query images by combining a support element and an image randomly sampled from the training dataset. This process ensures that the query contains sufficient information from the support image to be classified as the latter, while introducing diversity by blending them. Consequently, query images appear distinct from their corresponding support images while still belonging to the same class, better mimicking the tasks seen at test time and hence enhancing generalization. Following task creation, support and query images are encoded using a fixed pre-trained feature extractor. The resulting latent representations are aggregated into a sequence and passed as input to a transformer encoder along with their label encodings. The transformer encoder learns to extract contextual information from support images and predict the query image in a single pass, eliminating the need for the fine-tuning step during inference. An overview of our approach is visualized in Figure 1.
@@
Figure 1: Visualization of CAMeLU (example with 3-way 5-shot tasks). On the left side is the task creation mechanism, where $N$ samples are sampled from an unlabeled dataset $\mathcal{D}_{train}$. Each sample $x_{n}$ is augmented $K$ times to obtain $x_{n,k}$, and the same pseudo-label $n\in[1,N]$ is assigned to all data generated from the sample $x_{n}$. A strategy inspired by mixup [57] is utilized for generating the query set. On the right side, the so-created task is fed to the transformer encoder for predicting the query input. Inspired by CAML [12], we fed the transformer encoder with demonstrations built upon a concatenation of the features extracted from a fixed pre-trained feature extractor and a learned class encoder. The symbol * represents the unknown query label that the transformer encoder aims to predict.
@@
Throughout extensive experiments we demonstrate the effectiveness of the proposed approach to generalize to new tasks in real-time. Particularly, CAMeLU outperforms other UML baselines across all datasets, establishing itself as the state-of-the-art in the field. It also achieves comparable results to its supervised counterpart and to SSL approaches. While the latter requires fine-tuning on the test domain, CAMeLU obtains comparable performance with a single forward step, highlighting its applicability to real-time applications. Furthermore, by recasting the meta-learning phase as in-context learning within a transformer architecture, we improve efficiency, ensuring the whole training and inference phase can be executed with a consumer device with 4GB VRAM.
@@
  • We demonstrate that CAMeLU outperforms existing UML baselines across four datasets, without the need for fine-tuning to the test domains.
@@ 2 Related work
Meta-learning is a well-studied field in the machine learning community due to its ability to enable models to quickly adapt to tasks with limited labeled data. Pioneering work in the field [14, 40, 49, 32, 41] considers the scenarios where a large labeled dataset is available for meta-training, a challenging requirement in real-world applications. UML addresses this challenge by extracting meaningful information from unsupervised data that can be transferred to downstream tasks with limited labeled data. Different techniques have been explored in the literature to construct diverse tasks. CACTUs [18] applies clustering in the embedding space and assigns the same pseudo-label to all images in the same cluster. Other methods focus on generating synthetic samples, either using data augmentations, as in UMTRA [20], or leveraging interpolation in the latent space of a generative model [21]. Differently, Meta-GMVAE [25] and Meta-SVEBM [23] use variational autoencoders and memory-based models for pseudo-label generation. In recent methodologies, also SSL approaches [10] have been incorporated into UML methods. In particular, Set-SimCLR [24] builds on top of the SimCLR [7] approach and reframes meta-learning as a set-level problem, while PsCo [19], inspired by MoCo [16], utilizes a momentum encoder and a queue of previous samples to improve pseudo-labeling and construct diverse tasks for UML applications.
@@
Several UML approaches rely on data augmentation to construct the training tasks [20, 24, 19]. However, traditional transformations such as rotation, translation, cropping, resizing, and flipping [39] might generate images that are too similar to the original ones, ending up in tasks with low in-class variability between the support and query images. This creates a problem when the model needs to generalize to test tasks, where the query data are different instances than the support ones, not only augmented versions of them. In this paper, we addressed this limitation by generating query images using a strategy inspired by mixup [57] to enhance model generalization. Similarly to mixup [57], which performs linear interpolation of the feature vectors at the pixel level, other strategies based on mixing images comprise CutMix [56], PatchMix [28], and Manifold Mixup [47].
@@ 3.1 Task creation
$\textnormal{P}=1-\frac{(C\cdot m)\cdot((C-1)\cdot m)\cdots((C-N+1)\cdot m)}{(C\cdot m)\cdot(C\cdot m-1)\cdots(C\cdot m-N+1)}=\frac{C!\cdot m^{N}\cdot(C\cdot m-N)!}{(C-N)!\cdot(C\cdot m)!}.$
For example, the probability for a 5-way classification on the ImageNet-964 dataset used in our experiment is around $0.01$, which is negligible. To emulate the $K$-shot scenario typical of meta-learning tasks, we augment each of the $N$ images $K$ times, with an augmentation function $\mathcal{A}_{k}$ sampled from a predefined set of transformations $\mathbfcal{A}$, and we assign the same pseudo-label $n\in[1,N]$ to all data generated from the same sample $x_{n}$. Specifically, for each image $x_{n}$, $K$ augmentation functions are applied to obtain $x_{n,k}=\mathcal{A}_{k}(x_{n})$ with $\mathcal{A}_{k}\sim\mathbfcal{A}$ and $k=1,\ldots,K$. One requirement of $\mathcal{A}_{k}$ is that the function must preserve class membership, i.e., $x_{n}\in c\rightarrow\mathcal{A}_{k}(x_{n})\in c$, for $c\in C$. Although this property cannot be directly verified due to the lack of class information in the training set, it is reasonable to assume that it holds by selecting transformations that minimally alter the image content.
For the query set, we employ a different approach. We demonstrate in Appendix 7.2 that simply applying data augmentations sampled from $\mathbfcal{A}$ is not sufficient for creating a query set resembling those in test tasks. At test time, the query set samples are different instances belonging to the same $N$ classes encountered in the support set, not augmented versions of the support samples. However, since $\mathcal{D}_{train}$ is unlabeled, we need a strategy to create new samples with the same implicit classes as those in the support set. We propose a new strategy inspired by mixup [57]. For each query image $x_{j}^{(qr)}$ that we want to generate, we randomly sample an image $z_{j}$ from $\mathcal{D}_{train}$ and another image $x_{i}^{(sp)}$ from the support set. These images are then combined pixel-wise according to:
$x_{j}^{(qr)}=\lambda z_{j}+(1-\lambda)x_{i}^{(sp)}$ (1)
where $\lambda\sim Beta(\alpha,\beta)$ with $\alpha=1,\beta=1$ and $\lambda\in(0,0.5)$, and $x_{j}^{(qr)}$ is assigned the same label as $x_{i}^{(sp)}$. By merging a small proportion of a new image $z_{j}$ into $x_{i}^{(sp)}$, we enhance diversity in the query set with respect to the images in the support set. This strategy enforces the model to extract robust features and effectively generalize to scenarios where query images differ from the support samples, as commonly encountered at test time.
@@
While the strategy used for generating the query images draws inspiration from the mixup strategy proposed in [57], there are some substantial differences. The aim of mixup [57] is to develop a new data augmentation strategy to expand the number of training examples and diversify the data distribution used for training, thereby enhancing the robustness and generalization of neural networks. In CAMeLU, the primary objective of merging images is to encourage the model to learn even in scenarios where only a fraction of the class context is present in the image. In CAMeLU, $\lambda$ is sampled from a uniform distribution (obtained with a $Beta$ distribution with $\alpha=1,\beta=1$) in $(0,0.5)$, guaranteeing that the amount of information from $z_{j}$ that is embedded into $x_{j}^{(qr)}$ is less than 50%, which would otherwise corrupt the label data. Indeed, we assign the same label of $x_{i}^{(sp)}$ to $x_{j}^{(qr)}$, forcing the network to learn to retrieve information in $x_{j}^{(qr)}$ that is related to the category of $x_{i}^{(sp)}$. Contrarily, mixup creates new examples by interpolating both images and labels at the scope of extending the training distribution.
@@ 3.2 In-context learning method
To achieve this, we design a model comprising three components: (1) a feature extractor $f_{\psi}$, (2) a class encoder $g_{\phi}$, and (3) a transformer encoder with a linear projection layer on top, i.e., $M_{\theta}$. The feature extractor aims to map support and query samples into a latent space where images with similar characteristics and semantic meaning are assigned similar representations. Section 7.4 explores various feature extractors for this purpose, including those pre-trained via a supervised approach or leveraging an SSL technique. The resulting representations are then concatenated with a class embedding. The class embeddings for the support representations are generated by encoding the corresponding classes using the class encoder $g_{\phi}$. However, as the classes of the queries are unknown, a randomly initialized learnable vector is appended to each query representation. The so-combined embeddings are then organized into sequences $S_{i,j}=\left((f_{\psi}(x_{1}^{(sp)}),g_{\phi}(y_{1}^{(sp)})),\ldots,(f_{\psi}(x_{NK}^{(sp)}),g_{\phi}(y_{NK}^{(sp)})),f_{\psi}(x_{j}^{(qr)})\right),$ $j=1,\ldots,Q,$ resembling the one in Equation 2. These sequences are fed into the transformer encoder, and only the transformer output corresponding to the query sample is selected and passed through a projection layer to predict the query label. This process iterates for all queries in the task, and the aggregated loss is employed for model training. In particular, the training process can be formulated as an optimization process where the objective is as follows:
@@ 4 Experiments
In this section, we demonstrate the effectiveness of CAMeLU across different datasets, and we compare the results with several baseline methods. In particular, we provide a quantitative comparison with UML baselines in Sect. 4.3, and we highlight the ability of CAMeLU to leverage generalization over memorization in Sect. 4.4. We then present the results with a large-scale training dataset in Sect. 4.5 and a comparison with SSL methods in Sect. 4.6.
@@ 4.1 Datasets and baselines
For the evaluation, we use two generic object recognition datasets, i.e., miniImageNet [37] and CIFAR-fs [2], and two fine-grained image classification datasets, i.e., CUB [50], and Aircraft [30]. Each dataset is split into training, validation, and test sets following the splits in [37] and [2] for miniImageNet and CIFAR-fs, respectively, and in [43] for the remaining datasets. All labels are removed from the datasets during the training phase.
We compare the results of CAMeLU with standard UML approaches such as CACTUs [18], UMTRA [20], Meta-GMVAE [25], and PsCo [19]. These methods are evaluated in-domain as recommended in the original papers, with training and testing performed on the same dataset. While this setup is relatively simpler than the cross-domain evaluation employed for CAMeLU, applying these methods in a cross-domain scenario may not be fair, as they were not explicitly designed for such a challenging scenario. Only PsCo [19] is further evaluated in a cross-domain setting, as the authors demonstrate its adaptability to this scenario through an additional adaptation phase to the test domain. The performance of CAMeLU is also compared with CAML [12], a supervised meta-learning approach that assumes tasks are available both during the training and testing phases and leverages the in-context ability of transformer architectures to generalize to new tasks. Due to their similarities, we refer to CAML as the supervised counterpart of CAMeLU. Furthermore, Section 4.6 reports a comparative analysis with two fine-tuned state-of-the-art self-supervised trained networks, namely SimCLR [7] and SwAV [6].

### v2

@@ Abstract
Unsupervised meta-learning aims to learn feature representations from unsupervised datasets that can transfer to downstream tasks with limited labeled data. In this paper, we propose a novel approach to unsupervised meta-learning that leverages the generalization abilities of in-context learning observed in transformer architectures. Our method reframes meta-learning as a sequence modeling problem, enabling the transformer encoder to learn task context from support images and utilize it to predict query images. At the core of our approach lies the creation of diverse tasks generated using a combination of data augmentations and a mixing strategy that challenges the model during training while fostering generalization to unseen tasks at test time. Experimental results on benchmark datasets showcase the superiority of our approach over existing unsupervised meta-learning baselines, establishing it as the new state-of-the-art in the field. Remarkably, our method achieves competitive results with supervised and self-supervised approaches, underscoring the efficacy of the model in leveraging generalization over memorization.
@@ 1 Introduction
Various approaches have been proposed to address the UML problem [19, 20, 21, 24, 25, 26]. However, UML still faces several challenges. Existing UML methods often rely on simple data augmentations to construct the training tasks, while following the standard meta-learning task sampling pipeline for the evaluation phase. This results in a significant difference between training and testing tasks, constraining the model’s ability to generalize to unseen tasks during inference and requiring additional fine-tuning on the test domain. Furthermore, existing UML approaches typically assume that the training and test datasets belong to the same domain. In our framework, we loosen this assumption resulting in a more challenging setting that necessitates a better model generalization compared to usual meta-learning applications. We refer to this as the cross-domain scenario.
@@
In this paper, we propose a novel approach to UML that addresses these challenges by leveraging in-context learning within a transformer architecture [11, 32]. In-context learning allows the model to use the context provided by a sequence of input-output pairs to make predictions on new input data. Inspired by recent advancements in large language models (LLMs) [54, 4, 30], we formulate meta-learning as a sequence modeling problem, where a task is seen as a sequence of support images and an unknown query image. The support set is treated as the context utilized by the model to predict the class of the query image. We call our approach CAMeLU, which stands for Context-Aware Meta-Learning in Unsupervised scenarios. Central to our approach is a novel task creation mechanism that enables the generation of a large number of different tasks from an unlabeled dataset. Drawing inspiration from the natural decision process of learning by analogy [56], we construct tasks that closely resemble the structure of those encountered during inference. Specifically, we use a combination of different data augmentation techniques based on basic image manipulations [41] for generating the samples in the support set. Conversely, a strategy similar to mixup [60] is employed to generate query images by combining a support element, after applying a distinct augmentation function, and an image randomly sampled from the training dataset. This process ensures that the query contains sufficient information from the support image to be classified as the latter, while introducing diversity by blending them. Consequently, query images appear distinct from their corresponding support images while still belonging to the same class, better mimicking the tasks seen at test time and hence enhancing generalization. Following task creation, support and query images are encoded using a fixed pre-trained feature extractor. The resulting latent representations are aggregated into a sequence and passed as input to a transformer encoder along with their label encodings. The transformer encoder learns to extract contextual information from support images and predict the query image in a single pass, eliminating the need for the fine-tuning step during inference. An overview of our approach is visualized in Figure 1.
@@
Figure 1: Visualization of CAMeLU (with 3-way 5-shot tasks). The left side illustrates the task creation mechanism, where $N$ samples are drawn from an unlabeled dataset $\mathcal{D}_{train}$. Each sample $x_{n}$ is augmented $K$ times to obtain $x_{n,k}^{(sp)}$. A strategy inspired by mixup [60] is utilized for generating the query set by using an augmented version of $x_{n}$, i.e., $\tilde{x}_{n,j}$. The same pseudo-label $n\in[1,N]$ is assigned to all data generated from the sample $x_{n}$. On the right side, the so-created task is fed into the transformer encoder for predicting the query input. Inspired by CAML [12], the transformer encoder processes demonstrations created by concatenating features from a fixed pre-trained feature extractor and a learned class encoder. The symbol * denotes the unknown query label that the transformer encoder aims to predict.
@@
Throughout extensive experiments we demonstrate the effectiveness of the proposed approach to generalize to new tasks in real-time. Particularly, CAMeLU outperforms other UML baselines across several datasets, establishing itself as the state-of-the-art in the field. It also achieves comparable results to its supervised counterpart and to SSL approaches. While the latter requires fine-tuning on the test domain, CAMeLU obtains comparable performance with a single forward step, highlighting its applicability to real-time applications. Furthermore, by recasting the meta-learning phase as in-context learning within a transformer architecture, we improve efficiency, ensuring the whole training and inference phase can be executed with a consumer device with 8GB VRAM.
@@
  • We demonstrate that CAMeLU outperforms existing UML baselines across five datasets, without the need for fine-tuning to the test domains.
@@ 2 Related work
Meta-learning is a well-studied field in the machine learning community due to its ability to enable models to quickly adapt to tasks with limited labeled data. Pioneering work in the field [14, 42, 51, 33, 43] considers the scenarios where a large labeled dataset is available for meta-training, a challenging requirement in real-world applications. UML addresses this challenge by extracting meaningful information from unsupervised data that can be transferred to downstream tasks with limited labeled data. Different techniques have been explored in the literature to construct diverse tasks. CACTUs [19] applies clustering in the embedding space and assigns the same pseudo-label to all images in the same cluster. Other methods focus on generating synthetic samples, either using data augmentations, as in UMTRA [21], or leveraging interpolation in the latent space of a generative model [22]. Differently, Meta-GMVAE [26] and Meta-SVEBM [24] use variational autoencoders and memory-based models for pseudo-label generation. Recent methodologies have also incorporated SSL techniques [10] into UML methods. In particular, Set-SimCLR [25] builds on top of the SimCLR [7] approach and reframes meta-learning as a set-level problem, while PsCo [20], inspired by MoCo [16], utilizes a momentum encoder and a queue of previous samples to improve pseudo-labeling and construct diverse tasks for UML applications. Similarly, BECLR [34] introduces an approach for unsupervised few-shot learning by proposing a constrastive representation learning framework, instead of meta-learning.
@@
Several UML approaches rely on data augmentation to construct the training tasks [21, 25, 20]. However, traditional transformations such as rotation, translation, cropping, resizing, and flipping [41] might generate images that are too similar to the original ones, ending up in tasks with low in-class variability between the support and query images. This creates a problem when the model needs to generalize to test tasks, where the query data are different instances than the support ones, not only augmented versions of them. In this paper, we addressed this limitation by generating query images using a strategy inspired by mixup [60] to enhance model generalization. Similarly to mixup, which performs linear interpolation of the feature vectors at the pixel level, other strategies based on mixing images comprise CutMix [59], PatchMix [29], and Manifold Mixup [49].
@@ 3.1 Task creation
$\textnormal{P}=1-\frac{(C\cdot m)\cdot((C-1)\cdot m)\cdots((C-N+1)\cdot m)}{(C\cdot m)\cdot(C\cdot m-1)\cdots(C\cdot m-N+1)}=1-\frac{C!\cdot m^{N}\cdot(C\cdot m-N)!}{(C-N)!\cdot(C\cdot m)!}.$
For example, the probability for a 5-way classification on the ImageNet-964 dataset used in our experiment is around $0.01$, which is negligible. To emulate the $K$-shot scenario typical of meta-learning tasks, we augment each of the $N$ images $K$ times, with an augmentation function $\mathcal{A}_{k}$ sampled from a predefined set of transformations $\mathbfcal{A}$, and we assign the same pseudo-label $n\in[1,N]$ to all data generated from the same sample $x_{n}$. Specifically, for each image $x_{n}$, $K$ augmentation functions are applied to obtain $x_{n,k}^{(sp)}=\mathcal{A}_{k}(x_{n})$ with $\mathcal{A}_{k}\sim\mathbfcal{A}$ and $k=1,\ldots,K$. One requirement of $\mathcal{A}_{k}$ is that the function must preserve class membership, i.e., $x_{n}\in c\rightarrow\mathcal{A}_{k}(x_{n})\in c$, for $c\in C$. Although this property cannot be directly verified due to the lack of class information in the training set, it is reasonable to assume that it holds by selecting transformations that minimally alter the image content.
For the query set, we employ a different approach. We demonstrate in Appendix A.5 that simply applying data augmentations sampled from $\mathbfcal{A}$ is not sufficient for creating a query set resembling those in test tasks. At test time, the query set samples are different instances belonging to the same $N$ classes encountered in the support set, not augmented versions of the support samples. However, since $\mathcal{D}_{train}$ is unlabeled, we need a strategy to create new samples with the same implicit classes as those in the support set. For each query image $x_{j}^{(qr)}$ that we want to generate, we randomly select an image $x_{n}$ from the ones sampled for the support set generation and we apply an augmentation function $\mathcal{A}_{j}\sim\mathbfcal{A}$, possibly different from the one used for the support generation. We then propose a new strategy inspired by mixup, where we combined the augmented image $\tilde{x}_{n,j}=\mathcal{A}_{j}(x_{n})$ and an image $z_{j}$ sampled from $\mathcal{D}_{train}$ according to:
$x_{j}^{(qr)}=\lambda z_{j}+(1-\lambda)\tilde{x}_{n,j}$ (1)
where $\lambda\sim Beta(\alpha,\beta)$ with $\alpha=1,\beta=1$ and $\lambda\in(0,0.5)$, and $x_{j}^{(qr)}$ is assigned the same label $n$ as the support samples generated from $x_{n}$. By merging a small proportion of a new image $z_{j}$ into $\tilde{x}_{n,j}$, we enhance diversity in the query set with respect to the images in the support set. This strategy enforces the model to extract robust features and effectively generalize to scenarios where query images differ from the support samples, as commonly encountered at test time.
@@
While the strategy used for generating the query images draws inspiration from the mixup strategy proposed in [60], there are some substantial differences. The aim of mixup is to develop a new data augmentation strategy to expand the number of training examples and diversify the data distribution used for training, thereby enhancing the robustness and generalization of neural networks. In CAMeLU, the primary objective of merging images is to encourage the model to learn even in scenarios where only a fraction of the class context is present in the image. In CAMeLU, $\lambda$ is sampled from a uniform distribution (obtained with a $Beta$ distribution with $\alpha=1,\beta=1$) in $(0,0.5)$, guaranteeing that the amount of information from $z_{j}$ that is embedded into $x_{j}^{(qr)}$ is less than 50%, thus ensuring that the assigned label is consistent with the class of the support images generated from $x_{n}$. Indeed, we assign the same label $n$ to $x_{j}^{(qr)}$, forcing the network to learn to retrieve information in $x_{j}^{(qr)}$ that is related to the category of $x_{n}$. Contrarily, mixup creates new examples by interpolating both images and labels at the scope of limiting memorization over the training distribution.
@@ 3.2 In-context learning method
To achieve this, we design a model comprising three components: (1) a feature extractor $f_{\psi}$, (2) a class encoder $g_{\phi}$, and (3) a transformer encoder with a linear projection layer on top, i.e., $M_{\theta}$. The feature extractor aims to map support and query samples into a latent space where images with similar characteristics and semantic meaning are assigned similar representations. In Appendix A.4 we explore various feature extractors for this purpose, including those pre-trained via a supervised approach or leveraging an SSL technique. The resulting representations are then concatenated with a class embedding. The class embeddings for the support representations are generated by encoding the corresponding classes using the class encoder $g_{\phi}$. However, as the classes of the queries are unknown, a randomly initialized learnable vector is appended to each query representation. The so-combined embeddings are then organized into sequences $S_{i,j}=\left((f_{\psi}(x_{1}^{(sp)}),g_{\phi}(y_{1}^{(sp)})),\ldots,(f_{\psi}(x_{NK}^{(sp)}),g_{\phi}(y_{NK}^{(sp)})),f_{\psi}(x_{j}^{(qr)})\right),$ $j=1,\ldots,Q,$ resembling the one in Equation 2. These sequences are fed into the transformer encoder, and only the transformer output corresponding to the query sample is selected and passed through a projection layer to predict the query label. This process iterates for all queries in the task, and the aggregated loss is employed for model training. In particular, the training process can be formulated as an optimization process where the objective is as follows:
@@
During evaluation, when a new task is presented, the available examples in $D_{new}^{(sp)}$ are utilized as contextual information to guide the classfication of the query samples, without requiring any fine-tuning or adaptation steps.
@@ 4 Experiments
In this section, we demonstrate the effectiveness of CAMeLU across different datasets, and we compare the results with several baseline methods. In particular, we provide a quantitative comparison with UML baselines in Sect. 4.3, and we highlight the ability of CAMeLU to leverage generalization over memorization in Sect. 4.4. We then present the results with a small-scale training dataset in Sect. 4.5 and a comparison with SSL methods in Sect. 4.6.
@@ 4.1 Datasets and baselines
For the evaluation, we use two generic object recognition datasets, i.e., miniImageNet [39] and CIFAR-fs [2], and three fine-grained image classification datasets, i.e., CUB [52], Aircraft [31], and Meta-iNat [55]. While miniImageNet and CIFAR-fs share some similarities with ImageNet-1k, as some classes appears in all three datasets, CUB, Aircraft, and Meta-iNat focus on more specialized domains, ensuring a rigourous cross-domain evaluation. Each dataset is split into training, validation, and test sets following the splits in [39] and [2] for miniImageNet and CIFAR-fs, respectively, and in [45] and [34] for the remaining datasets. All labels are removed from the datasets during the training phase.
We compare the results of CAMeLU with standard UML approaches such as CACTUs [19], UMTRA [21], Meta-GMVAE [26], and PsCo [20]. These methods are evaluated in-domain as recommended in the original papers, with training and testing performed on the same dataset. While this setup is relatively simpler than the cross-domain evaluation employed for CAMeLU, applying these methods in a cross-domain scenario may not be fair, as they were not explicitly designed for such a challenging scenario. Only PsCo [20] is further evaluated in a cross-domain setting, as the authors demonstrate its adaptability to this scenario through an additional adaptation phase to the test domain. We also compare CAMeLU with BECLR [34] in Section 4.5, a contrastive framework for unsupervised few-shot learning. Although BECLR does not utilize meta-learning, it serves as a strong baseline for unsupervised few-shot learning. Additionally, CAMeLU is compared to CAML [12], a supervised meta-learning approach that assumes tasks are available both during the training and testing phases and leverages the in-context ability of transformer architectures to generalize to new tasks. Due to their similarities, we refer to CAML as the supervised counterpart of CAMeLU. Furthermore, Section 4.6 provides a comparative analysis with two fine-tuned state-of-the-art self-supervised trained networks, namely SimCLR [7] and SwAV [6].